# ContPhy: Continuum Physical Concept Learning and Reasoning from Videos

### v1

@@ 1 Introduction
As depicted in Fig. 1, objects like solids and liquids in nature often exhibit different properties, and these objects of different properties couple together to build our complex physical world. As humans, we are able to distinguish objects’ physical properties by observing their interactions. We know that the clear liquid in Fig. 1 (a) at the bottom has a higher density than the yellow liquid on the top; we know that the dynamic pulley in Fig. 1 (c) could help us to pull the cargo up more easily. These innate human skills raise an intriguing question: can current AI models have the physical common sense to infer physical properties of the continuum111Continuum encompasses various bodies like liquids, soft materials (e.g., ropes), rigid bodies and articulated bodies (e.g., pulleys). and predict corresponding dynamics?
@@
Recently, a series of benchmarks (Riochet et al., 2018; Rajani et al., 2020; Bear et al., ), have been developed to study machine models’ effectiveness for physical reasoning. However, there have been limitations that make them non-ideal for the assessment of whether machine models have human-like physical reasoning abilities. Firstly, most benchmarks mainly deal with simple visual primitives like spheres, cubes, and collision events of rigid objects only. It remains doubtful whether the conclusions based on these simple scenes will still hold in more comprehensive visual scenarios with the coupling of soft objects and their interaction with rigid objects. There have also been benchmarks like Physion (Bear et al., ) that were developed to evaluate machine models’ physical reasoning abilities in different scenarios. However, objects in Physion are of the same physical parameters without any variance (e.g. solids with the same mass and water with the same density). Moreover, Physion only requires models to predict whether two objects will come into contact after the observed video ends. It has not incorporated natural language to answer other challenging questions like predicting dynamics in counterfactual scenes and selecting actions to achieve a goal.
@@
To this end, we aim to build a Continuum Physical Dataset (ContPhy) to thoroughly evaluate and diagnose machine models’ physical reasoning performance in comprehensive physical environments. The design of ContPhy aims to achieve two goals: 1) covering diverse physical scenarios and 2) supporting comprehensive natural language tasks. To achieve the first goal, we adopt the physical engine (Haas, 2014) to simulate diverse videos with dense supervision signals. As shown in Fig. 2, the simulated physical scenes include scenes with the coupling of different liquids, deformable cloths, pulley systems, and elastoplastic balls. Another goal of the built dataset is to propose diverse physical reasoning tasks in the form of video question answering. We achieve this goal with a carefully designed question engine. The question engine takes the dense simulated video annotation as input and generates different questions based on pre-defined textual templates. Sample questions can be found in Fig. 2. It asks challenging questions such as “If the red stick were removed, would most orange fluid flow into the cyan container?” and “Is the mass of the sphere greater than half that of the red cube?”, which requires the model to have a deep understanding of physical scenes and reason about their dynamics.
@@
Dataset | Question | Rationales | Diverse | Goal-driven | Interaction | Counterfactual |
Answering | Scenarios | Questions | of soft objects | Property Dynamics | ||
IntPhys (Riochet et al., 2018) | ||||||
ESPRIT (Rajani et al., 2020) | ||||||
Cater (Girdhar & Ramanan, 2020) | ||||||
CoPhy(Baradel et al., 2020) | ||||||
CRAFT (Ates et al., 2020) | ||||||
CLEVRER (Yi et al., 2020) | ||||||
Physion (Bear et al., ) | ||||||
ComPhy (Chen et al., 2022) | ||||||
CRIPP-VQA (Patel et al., 2022) | ||||||
@@ 2 Related Work
Physical Reasoning. Our work is closely related to Physical Reasoning benchmarks (Rajani et al., 2020; Girdhar & Ramanan, 2020; Baradel et al., 2020; Bear et al., ; Li et al., 2022b, c). We summarize the key features of these various benchmarks and compare them against our benchmark in table 1. Early benchmarks (Riochet et al., 2018; Rajani et al., 2020) simulate physical scenes with visual primitives and test models’ physical intuition. Later, CLEVER (Yi et al., 2020), ComPhy (Chen et al., 2022), and CRIPP-VQA (Patel et al., 2022) extend the simple visual primitives with natural language and asked questions about rigid bodies’ collisions. Recently, Physion (Bear et al., ; Tung et al., 2023) provides more complex visual scenes and requires models to predict whether two objects will come into contact in future frames. As summarized in table 1, the proposed ContPhy is the only benchmark that contains soft objects with different physical parameters and asks diverse language-based questions about dynamics in counterfactual and goal-planning scenarios.
Video Question Answering. Our paper is also related to Video Question Answering (VQA) (Lei et al., 2018; Zadeh et al., 2019; Jang et al., 2017; Chen et al., 2021; Wu et al., 2021; Ding et al., 2021), which mainly requires machine models to answer questions about a given image or video’s content like visual attributes, actions, activity, and social events. However, existing VQA datasets (Zadeh et al., 2019; Xu et al., 2016; Wang et al., 2019) still typically assess abilities in visual perception, recognizing objects, shapes, and colors, and understanding human-centric actions. In this paper, we aim to build a benchmark that evaluates AI models’ comprehensive physical reasoning abilities.
Physical Benchmarks for Soft Bodies. Recently, there has been growing interest in the properties and dynamics of soft-bodied objects (Xiang et al., 2020; Gan et al., 2020; Macklin et al., 2014; Xian et al., 2023; Haas, 2014). Much of the research has concentrated on creating simulations of deformable objects and fluids to advance robotic manipulation and cognitive experimentation. Leveraging simulation tools, we can simulate deformable objects and fluids with varying physical parameters, enabling collaboration with natural language for physical commonsense reasoning.
@@ 3.1 Diverse Physical Scenarios
Diversity. The diversity of our video data arises from a wide range of materials with varying properties and changing dynamic phenomena. We include these materials and their physical properties in our scenarios compositionally, requiring models to obtain a deeper understanding of these scenes. First, this encompasses rigid bodies, ropes, cloths, plastoelastic balls, fluids, and articulated bodies, as well as their couplings, expanding our dataset diversity. For instance, Fig 1 (c) portrays the coupling of deformable ropes and articulated pulleys. Second, a key feature that distinguishes the proposed ContPhy dataset from existing benchmarks like CLEVRER (Yi et al., 2020) and Physion (Bear et al., ), is the inclusion of varying physical properties and commonsense concepts that are generally acknowledged by common people. These properties could scarcely be inferred through static images, involving mass, density, tension, friction, stretchiness, bending stiffness, elasticity, and plasticity. Such variation can alter the dynamics and generate differing future states, enriching the dataset with a wider range of observable physical behaviors. As an example, the liquid’s position in Fig. 1 (a) is jointly determined by its density and its interaction with the nearby liquids.
Scenarios. We design four physical dynamic scenarios to comprehensively benchmark models’ cognitive ability on the continuum. a) Liquid Dynamics: In Fig. 2 A, we present a liquid hourglass-like device. Different liquids, each with unique colors and densities, are released from upper emitters, flowing through fixed sticks, altering directions, and finally reaching containers at the bottom. This setup reveals distinct behaviors arising from interactions between fluids with varied densities and dynamic trajectories. Our research focuses on investigating the physical properties and trajectories of these liquids. b) Cloths Manipulation. As shown in Fig. 2 B, two cloth pieces with distinct stretching, bending, and frictional characteristics are pulled over objects, inducing potential collision events. The released fabric obstructs object views but outlines their shapes through deformations. Objects may topple if they surpass a height threshold or have low mass. This test assesses models’ ability to discern fabric properties and predict spatial behaviors of concealed objects based on the dynamic 3D surface geometry of the fabric. c) Rope Pulley System: In Fig. 2 C, a wall-mounted arrangement of pulleys, both movable and fixed and anchor points are depicted. Objects with varying masses interact, resulting in diverse motion patterns. The model’s main goal is to identify tension distributions within this basic rope system. It must also recognize correlations or constraints among moving objects, such as coordinated loads and pulley rotations on a single rope. Additionally, the model is expected to infer numerical relationships between loads’ masses and rope segment tensions. d) Soft Ball Dynamics: Fig. 2 D illustrates a playground with colored obstacles and randomly placed pits. Plastoelastic balls with different deformation resistance and yield stress are launched from varying positions, undergoing dynamic movements, including bouncing and permanent deformation. Some balls may collide with obstacles and fall into pits. This experiment assesses the model’s ability to accurately discern the elasticity and plasticity properties of soft bodies and predict their dynamic behavior.
@@ 3.2 Diverse Structured Questions
We categorize questions into two major groups: Physical Property Questions and Dynamics Questions. Fig. 2 shows the question types of the four scenarios. Sample templates are provided in the supplementary material.
Physical Property Questions. We formulated a set of physical property questions across four distinct scenarios. We pose questions about our chosen physical properties which can only be answered by observing object dynamics and interactions, neither static images nor single object behaviors. These questions can be answered with a brief phrase. Models are expected to deduce physical properties based on input video data, which requires physical common sense. Besides, we also inquire about the visible physical properties of objects, such as colors, shapes, and existences, which are shown in static frames.
Dynamics Questions. Dynamic questions can be further categorized into three types: counterfactual, goal-driven, and predictive, respectively concerning potential outcomes of changed conditions, strategies for specific objectives, and predictions about the future. In the fluid and ball scenarios, we crafted questions covering all three types, anticipating models to develop a comprehensive understanding of these scenarios through diverse question templates. For rope and cloth scenarios, we selectively assess a subset of dynamic question types due to scenario complexity. In the rope scenario, only counterfactual and goal-driven questions are included. In the cloth scenario, exclusively predictive questions prompt the model to anticipate outcomes not directly visible under the cloth cover. To increase cognitive challenge, we’ve designed multiple-choice questions with more than two but fewer than five answer choices, requiring models to provide a binary prediction for each option.
@@ 3.3 Generation Setup and Statistics
Video Generation. We used the Unity engine (Haas, 2014), an efficient platform, to simulate and render videos. We follow a bottom-up approach to generate videos and their annotation, involving the following sequential steps: a) Sampling: randomly select scene layouts, camera parameters, and initial conditions to create a diverse set of scenarios. b) Initialization: place and configure objects within the scene. c) Pre-simulation: conduct a preliminary simulation to evaluate whether the obtained simulation results align with the expected data distribution. d) Rendering: generate high-quality videos by configured cameras. e) Post-simulation: carry out multiple simulations under varying conditions and record the simulation outputs. f) Output: produce rich sensor data and annotation information, encompassing original video, segmentation, bounding boxes, particles, meshes, collision events, configurations, and other simulation raw data required for question generation. Supplementary material provides more details.
Question Generation. We develop a question engine to generate question-answering pairs step by step. a) Template Design: create massive question and option templates. b) Sampling: retrieve the simulation results, combine the properties of possible objects with predefined templates, sample questions, and options accordingly, and determine correct answers. Target objects possess unique names described by visual attributes like color, shape, orientation, and mobility. c) Re-Sampling: ensure a balanced distribution of answers among the options to prevent answer bias.
Question Statistics. The video content prompted the formulation of numerous questions, with each video featuring one property question and two dynamics questions, except for the rope scenario, which included two property-related questions and two dynamics questions. A total of 2,000 questions were generated for the rope scenario, and 1,500 questions were created for other scenarios. The dataset encompasses 6,500 questions derived from 2,000 videos. We divided the dataset into three subsets: 50% for training, 20% for validation, and 30% for testing. Across the entire dataset, 20% of questions are counterfactual, 11% are goal-driven, 22% are predictive, and the remaining 46% are related to various physical property questions. Further details on the distribution of each question type and templates for each scenario can be found in the supplementary material.
@@ 4 ContPRO
Inspired by prior work (Yi et al., 2018, 2020), we propose an oracle neural-symbolic framework named ContPRO for ContPhy. As shown in Fig. 3, we decompose the question-answering task into four main modules, video perception, physical simulation, program parser, and symbolic execution. Given a raw video, the video perception module detects the objects and their associated static attributes with MASK-RCNN detector (He et al., 2017). The physical simulator takes point clouds as input and predicts objects’ dynamics in different scenarios with dynamic prediction models (Jiang et al., 2015; Li et al., 2018). The program parser translates the question query into executable programs with a large language model (LLM). Based on the object attributes and dynamics, the symbolic executor executes the programs to get the answer to the question.
@@
Video Perception. The video perception module is supported by a MASK R-CNN (He et al., 2017) to densely detect objects’ location in each frame and associated static attributes like color and material. We take the ResNet-50 (He et al., 2016) as the backbone and fine-tune the network with data from the training set of all four scenarios until converges.
Physical Simulation. We choose DPI-Net (Li et al., 2018) for dynamic prediction for rope and cloth as it has shown reasonable dynamic prediction abilities in different materials (Chen et al., 2022; Bear et al., ). We also observe that the Material Point Method (MPM) (Sulsky et al., 1995) can estimate physical properties and dynamics better for fluids and objects of varying plasticity with its differentiable mechanism during inference. Thus, we adopt MPM to predict dynamics for scenes of fluid and soft balls.
@@
Large Language Model as Program Parser. Traditional neuro-symbolic models (Andreas et al., 2016; Yi et al., 2018) usually train a domain-specific sequence-to-sequence model to translate the natural language query into executable programs, which requires manual implementation of each symbolic operator and show problems in generalizing to questions out of the training set distribution. Motivated by the recent ViperGPT (Surís et al., 2023), we utilize the large language model, ChatGPT (Ouyang et al., 2022) as the language parser. We further develop a set of dynamics modules and visual perception modules serving as APIs for solution generation. With the provided API access and a pre-defined physical reasoning prompt, we leverage ChatGPT to generate Python code that can be directly executed and interpreted. This module bridges the gap between language comprehension and physical concept understanding.
Program Execution. After extracting objects’ static attributes, physical properties, and dynamic trajectories and parsing the natural language query into an executable program, we execute the program with object states as input and output the predicted answer. We provide more model details in the supplementary material.

### v2

@@ 1 Introduction
As depicted in Figure 1, objects like solids and liquids in nature often exhibit different properties, and these objects of different properties couple together to build our complex physical world. As humans, we are able to distinguish objects’ physical properties by observing their interactions.We know that the clear liquid in Figure 1 (a) at the bottom has a higher density than the yellow liquid on the top; we know that the dynamic pulley in Figure 1 (c) could help us to pull the cargo up more easily. These innate human skills raise an intriguing question: can current AI models have the physical common sense to infer physical properties of the continuum111Continuum encompasses various bodies like liquids, soft materials (e.g., ropes), rigid bodies and articulated bodies (e.g., pulleys). More details about the physical concept of the continuum can be found in Section 9 in the Appendix. and predict corresponding dynamics?
@@
Recently, a series of benchmarks (Riochet et al., 2018; Rajani et al., 2020; Bear et al., 2021), have been developed to study machine models’ effectiveness for physical reasoning. However, there have been limitations that make them non-ideal for the assessment of whether machine models have human-like physical reasoning abilities. Firstly, most benchmarks mainly deal with simple visual primitives like spheres, cubes, and collision events of rigid objects only. It remains doubtful whether the conclusions based on these simple scenes will still hold in more comprehensive visual scenarios with the coupling of soft objects and their interaction with rigid objects. There have also been benchmarks like Physion (Bear et al., 2021) that were developed to evaluate machine models’ physical reasoning abilities in different scenarios. However, objects in Physion are of the same physical parameters without any variance (e.g. solids with the same mass and water with the same density). Moreover, Physion only requires models to predict whether two objects will come into contact after the observed video ends. It has not incorporated natural language to answer other challenging questions like predicting dynamics in counterfactual scenes and selecting actions to achieve a goal.
To this end, we aim to build a Continuum Physical Dataset (ContPhy) to thoroughly evaluate and diagnose machine models’ physical reasoning performance in comprehensive physical environments. The design of ContPhy aims to achieve two goals: 1) covering diverse physical scenarios and 2) supporting comprehensive natural language tasks.
@@
To achieve the first goal, we adopt the physical engine (Haas, 2014) to simulate diverse videos with dense supervision signals. As shown in Figure 2, the simulated physical scenes include scenes with the coupling of different liquids, deformable cloths, pulley systems, and elastoplastic balls. Another goal of the built dataset is to propose diverse physical reasoning tasks in the form of video question answering. We achieve this goal with a carefully designed question engine. The question engine takes the dense simulated video annotation as input and generates different questions based on pre-defined textual templates. Sample questions can be found in Figure 2. It asks challenging questions such as “If the red stick were removed, would most orange fluid flow into the cyan container?” and “Is the mass of the sphere greater than half that of the red cube?”, which requires the model to have a deep understanding of physical scenes and reason about their dynamics.
@@
| Dataset | Question | Rationales | Diverse | Goal-driven | Interaction | Counterfactual | |
| Answering | Scenarios | Questions | of soft objects | Property Dynamics | |||
| IntPhys (Riochet et al., 2018) | |||||||
| ESPRIT (Rajani et al., 2020) | |||||||
| Cater (Girdhar & Ramanan, 2020) | |||||||
| CoPhy(Baradel et al., 2020) | |||||||
| CRAFT (Ates et al., 2020) | |||||||
| CLEVRER (Yi et al., 2020) | |||||||
| Physion (Bear et al., 2021) | |||||||
| ComPhy (Chen et al., 2022) | |||||||
| ACQUIRED (Wu et al., 2023) | |||||||
| CRIPP-VQA (Patel et al., 2022) | |||||||
@@ 2 Related Work
Physical Reasoning.
Our work is closely related to Physical Reasoning benchmarks (Rajani et al., 2020; Girdhar & Ramanan, 2020; Baradel et al., 2020; Bear et al., 2021; Li et al., 2022b, c). We summarize the key features of these various benchmarks and compare them against our benchmark in table 1. Early benchmarks (Riochet et al., 2018; Rajani et al., 2020) simulate physical scenes with visual primitives and test models’ physical intuition. Later, CLEVER (Yi et al., 2020), ComPhy (Chen et al., 2022), and CRIPP-VQA (Patel et al., 2022) extend the simple visual primitives with natural language and asked questions about rigid bodies’ collisions. Recently, Physion (Bear et al., 2021; Tung et al., 2023) provides more complex visual scenes and requires models to predict whether two objects will come into contact in future frames. As summarized in table 1, the proposed ContPhy is the only benchmark that contains soft objects with different physical parameters and asks diverse language-based questions about dynamics in counterfactual and goal-planning scenarios.
Video Question Answering.
Our paper is also related to Visual Question Answering (VQA) (Lei et al., 2018; Zadeh et al., 2019; Jang et al., 2017; Chen et al., 2021; Wu et al., 2021; Ding et al., 2021; Hong et al., 2023; Chen et al., 2024b, a), which mainly requires machine models to answer questions about a given image or video’s content like visual attributes, actions, activity, and social events. However, existing VQA datasets (Zadeh et al., 2019; Xu et al., 2016; Wang et al., 2019; Wu et al., 2023; Wang et al., 2024) still typically assess abilities in visual perception, recognizing objects, shapes, and colors, and understanding human-centric actions. In this paper, we aim to build a benchmark that evaluates AI models’ comprehensive physical reasoning abilities.
Physical Benchmarks for Soft Bodies.
Recently, there has been growing interest in the properties and dynamics of soft-bodied objects (Xiang et al., 2020; Gan et al., 2020; Macklin et al., 2014; Xian et al., 2023; Haas, 2014). Much of the research has concentrated on creating simulations of deformable objects and fluids to advance robotic manipulation and cognitive experimentation. Leveraging simulation tools, we can simulate deformable objects and fluids with varying physical parameters, enabling collaboration with natural language for physical commonsense reasoning.
@@ 3.1 Diverse Physical Scenarios
The diversity of our video data arises from a wide range of materials with varying properties and changing dynamic phenomena. We include these materials and their physical properties in our scenarios compositionally, requiring models to obtain a deeper understanding of these scenes. First, this encompasses rigid bodies, ropes, cloths, plastoelastic balls, fluids, and articulated bodies, as well as their couplings, expanding our dataset diversity. For instance, Figure 1 (c) portrays the coupling of deformable ropes and articulated pulleys. Second, a key feature that distinguishes the proposed ContPhy dataset from existing benchmarks like CLEVRER (Yi et al., 2020) and Physion (Bear et al., 2021), is the inclusion of varying physical properties and commonsense concepts that are generally acknowledged by common people. These properties could scarcely be inferred through static images, involving mass, density, tension, friction, stretchiness, bending stiffness, elasticity, and plasticity. Such variation can alter the dynamics and generate differing future states, enriching the dataset with a wider range of observable physical behaviors. As an example, the liquid’s position in Figure 1 (a) is jointly determined by its density and its interaction with the nearby liquids.
Scenarios.
We design four physical dynamic scenarios to comprehensively benchmark models’ cognitive ability on the continuum, as shown in Figure 2.
a) Liquid Dynamics. In Figure 2 A, we present a liquid hourglass-like device. Different liquids, each with unique colors and densities, are released from upper emitters, flowing through fixed sticks, altering directions, and finally reaching containers at the bottom. This setup reveals distinct behaviors arising from interactions between fluids with varied densities and dynamic trajectories. Our research focuses on investigating the physical properties and trajectories of these liquids.
b) Cloths Manipulation. As shown in Figure 2 B, two cloth pieces with distinct stretching, bending, and frictional characteristics are pulled over objects, inducing potential collision events. The released fabric obstructs object views but outlines their shapes through deformations. Objects may topple if they surpass a height threshold or have low mass. This test assesses models’ ability to discern fabric properties and predict spatial behaviors of concealed objects based on the dynamic 3D surface geometry of the fabric.
c) Rope Pulley System. In Figure 2 C, a wall-mounted arrangement of pulleys, both movable and fixed and anchor points are depicted. Objects with varying masses interact, resulting in diverse motion patterns. The model’s main goal is to identify tension distributions within this basic rope system. It must also recognize correlations or constraints among moving objects, such as coordinated loads and pulley rotations on a single rope. Additionally, the model is expected to infer numerical relationships between loads’ masses and rope segment tensions.
d) Soft Ball Dynamics. Figure 2 D illustrates a playground with colored obstacles and randomly placed pits. Plastoelastic balls with different deformation resistance and yield stress are launched from varying positions, undergoing dynamic movements, including bouncing and permanent deformation. Some balls may collide with obstacles and fall into pits. This experiment assesses the model’s ability to accurately discern the elasticity and plasticity properties of soft bodies and predict their dynamic behavior.
@@ 3.2 Diverse Structured Questions
We categorize questions into two major groups: Physical Property Questions and Dynamics Questions. Figure 2 shows the question types of the four scenarios. Sample templates are provided in Table 5, 6, 7, and 8 in the Appendix.
Physical Property Questions.
We formulated a set of physical property questions across four distinct scenarios. We pose questions about our chosen physical properties which can only be answered by observing object dynamics and interactions, neither static images nor single object behaviors. These questions can be answered with a brief phrase. Models are expected to deduce physical properties based on input video data, which requires physical common sense. Besides, we also inquire about the visible physical properties of objects, such as colors, shapes, and existences, which are shown in static frames.
Dynamics Questions.
Dynamic questions can be further categorized into three types: counterfactual, goal-driven, and predictive, respectively concerning potential outcomes of changed conditions, strategies for specific objectives, and predictions about the future. In the fluid and ball scenarios, we crafted questions covering all three types, anticipating models to develop a comprehensive understanding of these scenarios through diverse question templates. For rope and cloth scenarios, we selectively assess a subset of dynamic question types due to scenario complexity. In the rope scenario, only counterfactual and goal-driven questions are included. In the cloth scenario, exclusively predictive questions prompt the model to anticipate outcomes not directly visible under the cloth cover. To increase cognitive challenge, we’ve designed multiple-choice questions with more than two but fewer than five answer choices, requiring models to provide a binary prediction for each option.
@@ 3.3 Generation Setup and Statistics
Video Generation.
We used the Unity engine (Haas, 2014), an efficient platform, to simulate and render videos. We follow a bottom-up approach to generate videos and their annotation, involving the following sequential steps:
a) Sampling. Randomly select scene layouts, camera parameters, and initial conditions to create a diverse set of scenarios.
b) Initialization. Place and configure all objects within the scenarios.
c) Pre-simulation. Conduct a preliminary simulation to evaluate whether the obtained simulation results align with the expected data distribution.
d) Rendering. Generate high-quality videos with configured cameras.
e) Post-simulation. Carry out multiple simulations under varying conditions and record the simulation outputs.
f) Output. Produce rich sensor data and annotation information, encompassing original video, segmentation, bounding boxes, particles, meshes, collision events, configurations, and other simulation raw data required for question generation. We will provide more details in Section 8.1 in the Appendix.
Question Generation.
We develop a question engine to generate question-answering pairs step by step:
a) Template Design. Create massive question and option templates.
b) Sampling. Retrieve the simulation results, combine the properties of possible objects with predefined templates, sample questions, and options accordingly, and determine correct answers. Target objects possess unique names described by visual attributes like color, shape, orientation, and mobility.
c) Re-Sampling. Ensure a balanced distribution of answers among the options to prevent answer bias.
Question Statistics.
The video content prompted the formulation of numerous questions, with each video featuring one property question and two dynamics questions, except for the rope scenario, which included two property-related questions and two dynamics questions. A total of 2,000 questions were generated for the rope scenario, and 1,500 questions were created for other scenarios. The dataset encompasses 6,500 questions derived from 2,000 videos. We divided the dataset into three subsets: 50% for training, 20% for validation, and 30% for testing. Across the entire dataset, 20% of questions are counterfactual, 11% are goal-driven, 22% are predictive, and the remaining 46% are related to various physical property questions. Further details on the distribution of each question type and templates for each scenario can be found in Section 8.2 in the Appendix.
@@ 4 ContPRO
Inspired by prior work (Yi et al., 2018, 2020), we propose an oracle neural-symbolic framework named ContPRO for ContPhy. As shown in Figure 3, we decompose the question-answering task into four main modules, video perception, physical simulation, program parser, and symbolic execution. Given a raw video, the video perception module detects the objects and their associated static attributes with MASK-RCNN detector (He et al., 2017). The physical simulator takes point clouds as input and predicts objects’ dynamics in different scenarios with dynamic prediction models (Jiang et al., 2015; Li et al., 2018). The program parser translates the question query into executable programs with a large language model (LLM). Based on the object attributes and dynamics, the symbolic executor executes the programs to get the answer to the question.
@@
Video Perception.
The video perception module is supported by a MASK R-CNN (He et al., 2017) to densely detect objects’ location in each frame and associated static attributes like color and material. We take the ResNet-50 (He et al., 2016) as the backbone and fine-tune the network with data from the training set of all four scenarios until converges.
Physical Simulation.
We choose DPI-Net (Li et al., 2018) for dynamic prediction for rope and cloth as it has shown reasonable dynamic prediction abilities in different materials (Chen et al., 2022; Bear et al., 2021). We also observe that the Material Point Method (MPM) (Sulsky et al., 1995) can estimate physical properties and dynamics better for fluids and objects of varying plasticity with its differentiable mechanism during inference. Thus, we adopt MPM to predict dynamics for scenes of fluid and soft balls.
@@
Large Language Model as Program Parser.
Traditional neuro-symbolic models (Andreas et al., 2016; Yi et al., 2018) usually train a domain-specific sequence-to-sequence model to translate the natural language query into executable programs, which requires manual implementation of each symbolic operator and show problems in generalizing to questions out of the training set distribution. Motivated by the recent ViperGPT (Surís et al., 2023), we utilize the large language model, ChatGPT (Ouyang et al., 2022) as the language parser. We further develop a set of dynamics modules and visual perception modules serving as APIs for solution generation. With the provided API access and a pre-defined physical reasoning prompt, we leverage ChatGPT to generate Python code that can be directly executed and interpreted. This module bridges the gap between language comprehension and physical concept understanding.
Program Execution.
After extracting objects’ static attributes, physical properties, and dynamic trajectories and parsing the natural language query into an executable program, we execute the program with object states as input and output the predicted answer. We provide more model details in Section 10.2 in the Appendix.